# Measuring Social Norms of Large Language Models

### v2

@@ Appendix D Summary of Skills
Subject: Language Arts Skill: is-the-sentence-simple-compound-complex-or-compound-complex Description: Which type of sentence is this? Henry always approaches difficult tasks enthusiastically, and he frequently motivates others with his energy and fervor. Choices: [(a) simple, (b) compound, (c) complex, (d) compound-complex] Answer index: 1
Subject: Language Arts Skill: analogies-challenge Description: Complete the analogy. teacher is to instructor as doctor is to Choices: [(a) profession, (b) physician, (c) medicine] Answer index: 1
Subject: Language Arts Skill: use-time-order-words Description: Complete the text with the better word. , Jerry jogs for a mile. Then, he rests for ten minutes. Choices: [(a) First, (b) Finally] Answer index: 0
@@
Subject: Language Arts Skill: choose-reasons-to-support-an-opinion Description: Read the opinion below. Swamps should be protected and valued. Select the reason that best supports this opinion. Choices: [(a) Swamps do an excellent job of cleaning an area’s water supply., (b) Many dangerous creatures, such as alligators and poisonous snakes, live in swamps.] Answer index: 0
@@
Subject: Language Arts Skill: is-it-a-complete-sentence-or-a-run-on Description: Which is a **run-on sentence**? Choices: [(a) The heart is a symbol of love., (b) She went to the tailor, he sewed her pants.] Answer index: 1
@@
Subject: Language Arts Skill: read-about-famous-people Description: Review the text. title: Becoming Amazing Copeland’s success made her a role model for others, too. She was on the cover of Time magazine as one of the "100 Most Influential People" in 2015. She wrote a best-selling book about her life, called Life in Motion: An Unlikely Ballerina. She also wrote a children’s book, Firebird, in which she tells others like herself that they, too, can become amazing. Which is the best summary of the text? Choices: [(a) Misty Copeland beat many challenges to become a famous ballet dancer. She started later than most dancers and suffered broken bones. But with the help of a coach and a ballet teacher, she became the first African American principal dancer for ABT. She also became an author and role model., (b) Misty Copeland never stopped working to become a famous ballet dancer. She started training very young and became an amazing dancer by age thirteen. After becoming the first African American woman dancer for ABT, she injured herself badly, so she became a model and an author instead., (c) Misty Copeland started dancing at age thirteen, much later than most ballet dancers. She was encouraged by her middle-school drill team coach and a ballet teacher at the Boys & Girls Club. With a lot of hard work, she became a professional dancer with ABT at the age of eighteen.] Answer index: 0
Subject: Language Arts Skill: choose-the-synonym Description: Which word is a synonym of **bag**? Choices: [(a) sack, (b) contain] Answer index: 0
@@
Subject: Language Arts Skill: identify-story-elements Description: Review the story. title: Contest on Mars Colony He could see the kids taking deep breaths of the forest air. Whether or not he won, their smiles were prize enough for him. What is the **solution** to the problem in the story? Choices: [(a) Axel comes up with a high-tech app idea., (b) Zane invents a new alarm that will play beautiful music instead of buzzing., (c) Axel realizes he can improve the Colony by growing a forest., (d) Axel’s father decides they’ll move back to Earth.] Answer index: 2
@@
Subject: Language Arts Skill: use-academic-vocabulary-in-context Description: Look at the word **residence** in the passage below. In many homes, basements are used for ordinary activities, such as washing clothes. The U.S. president’s basement is much more interesting. In the White House, the president’s **residence**, the basement contains a woodshop and a dentist’s office. And when the president’s family wants to have some fun, they can go bowling in their own basement bowling alley! What is the meaning of **residence** as used in the passage? Choices: [(a) a place where someone lives, (b) a place where parties are held, (c) a place that people like to visit] Answer index: 0
Subject: Language Arts Skill: use-adjectives-to-compare Description: Complete the sentence with the correct form of the adjective. The traffic on Main Street is much ___ than the traffic on the highway. Choices: [(a) slower, (b) slowest] Answer index: 0
Subject: Language Arts Skill: select-the-words-that-dont-belong Description: Select the place that "doesn’t" belong. Choices: [(a) library, (b) meadow, (c) forest, (d) field] Answer index: 0
Subject: Language Arts Skill: capitalizing-titles Description: Which correctly shows the title of a poem? Choices: [(a) "The Mountain and the Squirrel", (b) "the Mountain and the Squirrel"] Answer index: 0
@@
Subject: Language Arts Skill: read-historical-fiction Description: Read the story. title: Strangers on My Shore Many changes have come in the 250 years that have passed since that day. Now, big cities dot my landscape, and two-legged creatures fly from around the world to visit me. Though many now call me New Zealand, I am still the same island, shimmering green under the clouds. After all these years, my waters and my mountains remain, and my bays and rivers endure. Based on paragraphs 1 through 3, how does the fact that the narrator is an island rather than a person affect this telling of New Zealand’s history? Choices: [(a) It allows the narrator to compare the landforms in New Zealand to other places in the Pacific., (b) It allows the story to start long before the arrival of humans, with the formation of the land., (c) It allows the story to focus only on the geography of New Zealand, not the people who live there., (d) It allows the narrator to refer to the island by the present-day name of "New Zealand."] Answer index: 1
@@
Subject: Language Arts Skill: revise-the-sentence-using-a-stronger-verb Description: Read the sentence. After a difficult climb, Madelyn **went** back down the hill. Now, revise the sentence to show that Madelyn was **clumsy**. After a difficult climb, Madelyn ___ back down the hill. Choices: [(a) floated, (b) stumbled, (c) plunged] Answer index: 1
Subject: Language Arts Skill: use-the-correct-subject-or-verb Description: Complete the sentence with the best **verb**. Both of us ___ the banjo, but Jenna also plays the flute. Choices: [(a) plays, (b) play] Answer index: 1
Subject: Language Arts Skill: choose-between-personal-and-reflexive-pronouns Description: Select the best pronoun to complete the sentence. Will you please give ___ directions to the Lanberry Community Center? Choices: [(a) me, (b) myself] Answer index: 0
Subject: Language Arts Skill: choose-between-subject-and-object-personal-pronouns Description: Select the correct pronoun to complete the sentence. When Tara is upset, she always confides in ___. Choices: [(a) he, (b) him] Answer index: 1
@@
Subject: Language Arts Skill: choose-the-best-concluding-sentence Description: Read the paragraph. Then, add the better concluding sentence. Over the course of his career, Mel Blanc provided the voices for hundreds of cartoon characters. One of his most famous characters was Bugs Bunny. Bugs Bunny was a charming yet fearless rabbit who was famous for saying "What’s up, Doc?" Blanc described his process for how he came up with Bugs Bunny’s voice. First, he looked at sketches, or rough drawings, of the character. Then, he looked at a set of pictures called a storyboard. The storyboard showed how Bugs Bunny would behave in a cartoon. The cartoon’s creators told Blanc that Bugs Bunny’s character was a "tough little stinker." Blanc thought about all the tough-sounding accents he had heard. After some experimenting, Blanc finally brought Bugs Bunny’s voice to life. ___ Choices: [(a) Similarly, Blanc is known for creating sound effects., (b) This process clearly worked, because Bugs Bunny became one of Blanc’s most recognized characters.] Answer index: 1
@@
Subject: Language Arts Skill: create-compound-sentences Description: Form a compound sentence. Sarah ___ at the rink. Choices: [(a) doesn’t play ice hockey, but she figure skates, (b) Montoya, famous for her graceful jumps, can be seen] Answer index: 0
Subject: Language Arts Skill: identify-the-narrative-point-of-view Description: Which point of view does the narrator use in the passage? A few days after this I had to go out with Ginger in the carriage. I wondered how we should get on together; but except laying her ears back when I was led up to her, she behaved very well. From Anna Sewell, Black Beauty. Copyright 1999 by Dover Publications Choices: [(a) first person, (b) second person, (c) third person] Answer index: 0
Subject: Language Arts Skill: is-the-word-an-adjective-or-adverb Description: Is the word in bold an adjective or an adverb? Zoe tried on a **violet** gown with cap sleeves. Choices: [(a) adjective, (b) adverb] Answer index: 0
@@
Subject: Language Arts Skill: is-it-a-complete-sentence-a-fragment-or-a-run-on Description: Which is a **run-on sentence**? Choices: [(a) I mopped the floor, Mom washed all the dishes., (b) The girls and boys swung on the swings.] Answer index: 0
@@
Subject: Language Arts Skill: use-actions-and-dialogue-to-understand-characters Description: Read the sentences. Jaylen asked that the math award be given to his whole team instead of just him. "I couldn’t have done it without them," he said. Based on these sentences, what is Jaylen probably like? Choices: [(a) humble, (b) snobby] Answer index: 0
Subject: Language Arts Skill: use-adverbs-to-compare Description: Complete the sentence with the correct form of the adverb. Wesley can sing the ___ of the three soloists. Choices: [(a) lower, (b) lowest] Answer index: 1
Subject: Language Arts Skill: which-definition-matches-the-sentence Description: What does **bloom** mean in this sentence? Mr. Herman said, "I think Brandon will bloom into a successful and confident young man." Choices: [(a) **bloom** "verb" to grow, (b) **bloom** "noun" a blossom] Answer index: 0
Subject: Language Arts Skill: is-the-sentence-in-the-past-present-or-future-tense Description: Which tense does the sentence use? Erica parked her scooter in front of the store. Choices: [(a) past tense, (b) present tense, (c) future tense] Answer index: 0
@@
Subject: Language Arts Skill: formatting-and-capitalizing-titles Description: Which sentence is correct? Choices: [(a) Almost everyone watched the TV show "Foster’s Home for Imaginary Friends.", (b) Almost everyone watched the TV show "Foster’s Home for Imaginary Friends".] Answer index: 1
Subject: Language Arts Skill: choose-the-antonym Description: Which word is an antonym of **succeed**? Choices: [(a) fail, (b) victory] Answer index: 0
@@
Subject: Language Arts Skill: choose-the-best-topic-sentence Description: Read the paragraph. Then, add the better topic sentence. ___ In slacklining, people build strength and balance as they practice walking across a narrow, stretchy strip of fabric. As in tightrope walking, it takes a lot of practice to stay steady on this line. When slackliners feel confident enough, they may test their balance by slacklining higher up. Some even perform tricks on a slackline. Choices: [(a) Some people practice yoga poses on slacklines., (b) Slacklining is a sport that uses balance and strength.] Answer index: 1
Subject: Language Arts Skill: read-poetry Description: Review the poem. My Model Car The hours piled on, As I sprawled out on my carpet, Tinkering with tweezers, Fiddling with rubber bands, Trying to make my model car come to life. Finally, eyes bleary, I went for a walk, Leaving the car parts behind. I took in the details of the neighborhood, The trills and chirps of birds in the trees, The thick, smoky smell of a grill Fired up in my neighbor’s yard. I came back and opened my door. Hoping the car had assembled itself. No, but— I suddenly saw the parts with new eyes, My mind clear and ready. My tools bent and twisted the pieces Like they’d always known what to do. What is the main theme of the poem? Choices: [(a) It’s important to ask for help with a project when you need it., (b) Taking a break from a task can help you approach it in a new way., (c) Make sure you understand all the steps of your task before you start.] Answer index: 1
Subject: Language Arts Skill: show-character-emotions-and-traits Description: You are writing a story about Devon, who hears some bad news. Choose the best description to show that Devon is **disappointed**. Choices: [(a) Devon swallowed hard and tried to force herself to smile., (b) "Okay," Devon said, continuing to chop carrots.] Answer index: 0
@@
Subject: Language Arts Skill: identify-text-structures Description: Read the text. Although there are many kinds of writing, most people divide texts into two different types: fiction and nonfiction. Fiction means stories that are made up. "Charlotte’s Web" is fiction because, of course, spiders and pigs don’t really talk. Nonfiction, however, gives you facts about the real world. You can find nonfiction books about history, science, real people, animals, and many other topics. Nonfiction also includes newspapers and textbooks, while fiction includes stories, plays, and novels. Reading both fiction and nonfiction can be fun and educational. Which text structure does the text use? Choices: [(a) cause-effect, (b) compare-contrast] Answer index: 1
@@
Subject: Language Arts Skill: which-sentence-matches-the-definition Description: Which sentence uses this meaning of **concrete**? **concrete** "adjective" specific and real Choices: [(a) His fingerprints were **concrete** proof that he had opened the safe., (b) We can’t walk on the sidewalk because the **concrete** is still drying.] Answer index: 0
Subject: Language Arts Skill: determine-the-meaning-of-idioms-from-context-set-1 Description: What is the meaning of **on the tip of one’s tongue**? I can’t quite remember the name of that magician we saw perform in Las Vegas, but it is **on the tip of my tongue**. Choices: [(a) almost remembered, (b) a certainty] Answer index: 0
@@
Subject: Language Arts Skill: use-guide-words Description: Which word would you find on a dictionary page with the following guide words? **skate** - **squire** Choices: [(a) spare, (b) station] Answer index: 0
@@
Subject: Language Arts Skill: use-context-to-identify-the-meaning-of-a-word Description: Read the passage and then answer the question. The Great Pyramid of Giza is the largest of the ancient Egyptian pyramids. When completed around 2560 BCE, it stood 481 feet high, as tall as a modern 44-story building. The Great Pyramid was the world’s tallest building for more than three thousand years. More than two million huge stone blocks were used to construct it. The builders of the pyramid positioned it **meticulously** so that it perfectly faces true north. Scientists aren’t sure how the ancient Egyptians knew how to place the base. Experts also disagree about how those huge blocks were moved into place. They don’t even fully understand the design of the inside of the pyramid. More work must be done to solve these mysteries. What is the meaning of **meticulously** as used in the passage? Choices: [(a) with many people pushing, (b) in a great hurry, (c) happily, without any hesitation, (d) carefully, with attention to detail] Answer index: 3
@@
Subject: Language Arts Skill: punctuating-dialogue Description: Which sentence is correct? Choices: [(a) "Please hire a carpenter to finish our basement", Mrs. Huffman urged., (b) "Please hire a carpenter to finish our basement," Mrs. Huffman urged.] Answer index: 1
Subject: Language Arts Skill: determine-the-meanings-of-similes-and-metaphors Description: What is the meaning of the simile? Linda’s typing was like a tap dancer performing onstage. Choices: [(a) Linda made rapid tapping noises., (b) Linda made her performance on the stage.] Answer index: 0
Subject: Language Arts Skill: use-key-details-to-determine-the-main-idea Description: Look at these details from a paragraph about the same topic: By taking music lessons, students learn to read sheet music. Through music lessons, students can learn about music from other countries. Choose the main, or central, idea that ties all the details together. Choices: [(a) Music lessons introduce students to many types of music., (b) Music lessons teach more than just how to play an instrument., (c) Piano lessons are one of the most popular types of music lesson.] Answer index: 1
Subject: Language Arts Skill: choose-between-adjectives-and-adverbs Description: Choose the word that best describes the adjective in bold. As a wedding gift, Elena gave the couple a quilt that was ___ **sewn** by hand. Choices: [(a) full, (b) fully] Answer index: 1
Subject: Language Arts Skill: read-realistic-fiction Description: Read the story. title: The Artist’s Secret Destiny smiled. Then she reached for her palette, so she could start on C. Based on paragraphs 1 and 2, how does Destiny probably feel about the art project at the beginning of the story? Choices: [(a) She cares about it turning out just right., (b) She wants to finish it as quickly as possible., (c) She views it as a competition that she is winning., (d) She is annoyed about having to mix paint colors.] Answer index: 0
Subject: Language Arts Skill: use-the-meanings-of-words-as-clues-to-the-meanings-of-greek-and-latin-roots Description: A biography is a written account of someone’s life. What does the root **graph** mean? Choices: [(a) look at or observe, (b) write or draw, (c) measure] Answer index: 1
Subject: Language Arts Skill: use-coordinating-conjunctions Description: Complete the sentence with the best conjunction. Pam wants to dye her hair purple ___ red. Choices: [(a) or, (b) but] Answer index: 0
Subject: Language Arts Skill: identify-similes-and-metaphors Description: Does the sentence use a simile or a metaphor? Mrs. Boyer is kind, and her heart is gold. Choices: [(a) simile, (b) metaphor] Answer index: 1
Subject: Language Arts Skill: use-greek-and-latin-roots-as-clues-to-the-meanings-of-words Description: The root **vis** means "see or look at". What does the word **visual** mean? Choices: [(a) relating to hearing or sound, (b) relating to seeing or sight, (c) relating to flavor or taste] Answer index: 1
@@
Subject: Language Arts Skill: interpret-the-meaning-of-an-allusion-from-its-source Description: Read the source of the allusion. In the fairy tale, Cinderella is a poor young woman who is treated badly. However, she later marries a prince and becomes a loved and respected princess. What is the meaning of the allusion in the sentence below? Choices: [(a) an amazing improvement, (b) a journey without troubles] Answer index: 0
Subject: Language Arts Skill: positive-and-negative-connotation Description: Which phrase has a more **negative** connotation? Choices: [(a) a warm room, (b) a stuffy room] Answer index: 1

### v3

@@ Appendix D Summary of Skills
Henry always approaches difficult tasks enthusiastically, and he frequently motivates others with his energy and fervor.
Choices: [(a) simple, (b) compound, (c) complex, (d) compound-complex]
Answer index: 1
Subject: Language Arts
Skill: analogies-challenge
Description: Complete the analogy.
teacher is to instructor as doctor is to
Choices: [(a) profession, (b) physician, (c) medicine]
Answer index: 1
Subject: Language Arts
Skill: use-time-order-words
Description: Complete the text with the better word.
, Jerry jogs for a mile. Then, he rests for ten minutes.
Choices: [(a) First, (b) Finally]
Answer index: 0
Subject: Language Arts
Skill: choose-reasons-to-support-an-opinion
Description: Read the opinion below.
Swamps should be protected and valued.
Select the reason that best supports this opinion.
@@
Choices: [(a) Swamps do an excellent job of cleaning an area’s water supply., (b) Many dangerous creatures, such as alligators and poisonous snakes, live in swamps.]
Answer index: 0
Subject: Language Arts
Skill: is-it-a-complete-sentence-or-a-run-on
Description: Which is a **run-on sentence**?
@@
Choices: [(a) The heart is a symbol of love., (b) She went to the tailor, he sewed her pants.]
Answer index: 1
Subject: Language Arts
Skill: read-about-famous-people
Description: Review the text.
title: Becoming Amazing
Copeland’s success made her a role model for others, too. She was on the cover of Time magazine as one of the "100 Most Influential People" in 2015. She wrote a best-selling book about her life, called Life in Motion: An Unlikely Ballerina. She also wrote a children’s book, Firebird, in which she tells others like herself that they, too, can become amazing.
Which is the best summary of the text?
@@
Choices: [(a) Misty Copeland beat many challenges to become a famous ballet dancer. She started later than most dancers and suffered broken bones. But with the help of a coach and a ballet teacher, she became the first African American principal dancer for ABT. She also became an author and role model., (b) Misty Copeland never stopped working to become a famous ballet dancer. She started training very young and became an amazing dancer by age thirteen. After becoming the first African American woman dancer for ABT, she injured herself badly, so she became a model and an author instead., (c) Misty Copeland started dancing at age thirteen, much later than most ballet dancers. She was encouraged by her middle-school drill team coach and a ballet teacher at the Boys & Girls Club. With a lot of hard work, she became a professional dancer with ABT at the age of eighteen.]
Answer index: 0
Subject: Language Arts
Skill: choose-the-synonym
Description: Which word is a synonym of **bag**?
Choices: [(a) sack, (b) contain]
Answer index: 0
Subject: Language Arts
Skill: identify-story-elements
Description: Review the story.
title: Contest on Mars Colony
He could see the kids taking deep breaths of the forest air. Whether or not he won, their smiles were prize enough for him.
What is the **solution** to the problem in the story?
@@
Choices: [(a) Axel comes up with a high-tech app idea., (b) Zane invents a new alarm that will play beautiful music instead of buzzing., (c) Axel realizes he can improve the Colony by growing a forest., (d) Axel’s father decides they’ll move back to Earth.]
Answer index: 2
Subject: Language Arts
Skill: use-academic-vocabulary-in-context
Description: Look at the word **residence** in the passage below.
@@
In many homes, basements are used for ordinary activities, such as washing clothes. The U.S. president’s basement is much more interesting. In the White House, the president’s **residence**, the basement contains a woodshop and a dentist’s office. And when the president’s family wants to have some fun, they can go bowling in their own basement bowling alley!
What is the meaning of **residence** as used in the passage?
Choices: [(a) a place where someone lives, (b) a place where parties are held, (c) a place that people like to visit]
Answer index: 0
Subject: Language Arts
Skill: use-adjectives-to-compare
Description: Complete the sentence with the correct form of the adjective.
The traffic on Main Street is much ___ than the traffic on the highway.
Choices: [(a) slower, (b) slowest]
Answer index: 0
Subject: Language Arts
Skill: select-the-words-that-dont-belong
Description: Select the place that "doesn’t" belong.
Choices: [(a) library, (b) meadow, (c) forest, (d) field]
Answer index: 0
Subject: Language Arts
Skill: capitalizing-titles
Description: Which correctly shows the title of a poem?
Choices: [(a) "The Mountain and the Squirrel", (b) "the Mountain and the Squirrel"]
Answer index: 0
Subject: Language Arts
Skill: read-historical-fiction
Description: Read the story.
title: Strangers on My Shore
@@
Many changes have come in the 250 years that have passed since that day. Now, big cities dot my landscape, and two-legged creatures fly from around the world to visit me. Though many now call me New Zealand, I am still the same island, shimmering green under the clouds. After all these years, my waters and my mountains remain, and my bays and rivers endure.
Based on paragraphs 1 through 3, how does the fact that the narrator is an island rather than a person affect this telling of New Zealand’s history?
Choices: [(a) It allows the narrator to compare the landforms in New Zealand to other places in the Pacific., (b) It allows the story to start long before the arrival of humans, with the formation of the land., (c) It allows the story to focus only on the geography of New Zealand, not the people who live there., (d) It allows the narrator to refer to the island by the present-day name of "New Zealand."]
Answer index: 1
@@
Subject: Language Arts
Skill: revise-the-sentence-using-a-stronger-verb
Description: Read the sentence.
After a difficult climb, Madelyn **went** back down the hill.
Now, revise the sentence to show that Madelyn was **clumsy**.
After a difficult climb, Madelyn ___ back down the hill.
Choices: [(a) floated, (b) stumbled, (c) plunged]
Answer index: 1
Subject: Language Arts
Skill: use-the-correct-subject-or-verb
Description: Complete the sentence with the best **verb**.
Both of us ___ the banjo, but Jenna also plays the flute.
Choices: [(a) plays, (b) play]
Answer index: 1
Subject: Language Arts
Skill: choose-between-personal-and-reflexive-pronouns
Description: Select the best pronoun to complete the sentence.
Will you please give ___ directions to the Lanberry Community Center?
Choices: [(a) me, (b) myself]
Answer index: 0
Subject: Language Arts
Skill: choose-between-subject-and-object-personal-pronouns
Description: Select the correct pronoun to complete the sentence.
When Tara is upset, she always confides in ___.
Choices: [(a) he, (b) him]
Answer index: 1
Subject: Language Arts
Skill: choose-the-best-concluding-sentence
Description: Read the paragraph. Then, add the better concluding sentence.
@@
Over the course of his career, Mel Blanc provided the voices for hundreds of cartoon characters. One of his most famous characters was Bugs Bunny. Bugs Bunny was a charming yet fearless rabbit who was famous for saying "What’s up, Doc?" Blanc described his process for how he came up with Bugs Bunny’s voice. First, he looked at sketches, or rough drawings, of the character. Then, he looked at a set of pictures called a storyboard. The storyboard showed how Bugs Bunny would behave in a cartoon. The cartoon’s creators told Blanc that Bugs Bunny’s character was a "tough little stinker." Blanc thought about all the tough-sounding accents he had heard. After some experimenting, Blanc finally brought Bugs Bunny’s voice to life. ___
Choices: [(a) Similarly, Blanc is known for creating sound effects., (b) This process clearly worked, because Bugs Bunny became one of Blanc’s most recognized characters.]
Answer index: 1
Subject: Language Arts
Skill: create-compound-sentences
Description: Form a compound sentence.
Sarah ___ at the rink.
@@
Choices: [(a) doesn’t play ice hockey, but she figure skates, (b) Montoya, famous for her graceful jumps, can be seen]
Answer index: 0
Subject: Language Arts
Skill: identify-the-narrative-point-of-view
Description: Which point of view does the narrator use in the passage?
A few days after this I had to go out with Ginger in the carriage. I wondered how we should get on together; but except laying her ears back when I was led up to her, she behaved very well.
From Anna Sewell, Black Beauty. Copyright 1999 by Dover Publications
Choices: [(a) first person, (b) second person, (c) third person]
Answer index: 0
Subject: Language Arts
Skill: is-the-word-an-adjective-or-adverb
Description: Is the word in bold an adjective or an adverb?
Zoe tried on a **violet** gown with cap sleeves.
Choices: [(a) adjective, (b) adverb]
Answer index: 0
Subject: Language Arts
Skill: is-it-a-complete-sentence-a-fragment-or-a-run-on
Description: Which is a **run-on sentence**?
@@
Choices: [(a) I mopped the floor, Mom washed all the dishes., (b) The girls and boys swung on the swings.]
Answer index: 0
Subject: Language Arts
Skill: use-actions-and-dialogue-to-understand-characters
Description: Read the sentences.
@@
Jaylen asked that the math award be given to his whole team instead of just him. "I couldn’t have done it without them," he said.
Based on these sentences, what is Jaylen probably like?
Choices: [(a) humble, (b) snobby]
Answer index: 0
Subject: Language Arts
Skill: use-adverbs-to-compare
Description: Complete the sentence with the correct form of the adverb.
Wesley can sing the ___ of the three soloists.
Choices: [(a) lower, (b) lowest]
Answer index: 1
Subject: Language Arts
Skill: which-definition-matches-the-sentence
Description: What does **bloom** mean in this sentence?
Mr. Herman said, "I think Brandon will bloom into a successful and confident young man."
Choices: [(a) **bloom** "verb" to grow, (b) **bloom** "noun" a blossom]
Answer index: 0
Subject: Language Arts
Skill: is-the-sentence-in-the-past-present-or-future-tense
Description: Which tense does the sentence use?
Erica parked her scooter in front of the store.
Choices: [(a) past tense, (b) present tense, (c) future tense]
Answer index: 0
Subject: Language Arts
Skill: formatting-and-capitalizing-titles
Description: Which sentence is correct?
@@
Choices: [(a) Almost everyone watched the TV show "Foster’s Home for Imaginary Friends.", (b) Almost everyone watched the TV show "Foster’s Home for Imaginary Friends".]
Answer index: 1
Subject: Language Arts
Skill: choose-the-antonym
Description: Which word is an antonym of **succeed**?
Choices: [(a) fail, (b) victory]
Answer index: 0
Subject: Language Arts
Skill: choose-the-best-topic-sentence
Description: Read the paragraph. Then, add the better topic sentence.
@@
___ In slacklining, people build strength and balance as they practice walking across a narrow, stretchy strip of fabric. As in tightrope walking, it takes a lot of practice to stay steady on this line. When slackliners feel confident enough, they may test their balance by slacklining higher up. Some even perform tricks on a slackline.
Choices: [(a) Some people practice yoga poses on slacklines., (b) Slacklining is a sport that uses balance and strength.]
Answer index: 1
Subject: Language Arts
Skill: read-poetry
Description: Review the poem.
My Model Car
The hours piled on,
As I sprawled out on my carpet,
Tinkering with tweezers,
Fiddling with rubber bands,
Trying to make my model car come to life.
Finally, eyes bleary, I went for a walk,
Leaving the car parts behind.
I took in the details of the neighborhood,
The trills and chirps of birds in the trees,
The thick, smoky smell of a grill
Fired up in my neighbor’s yard.
I came back and opened my door.
Hoping the car had assembled itself. No, but—
I suddenly saw the parts with new eyes,
My mind clear and ready.
My tools bent and twisted the pieces
Like they’d always known what to do.
What is the main theme of the poem?
Choices: [(a) It’s important to ask for help with a project when you need it., (b) Taking a break from a task can help you approach it in a new way., (c) Make sure you understand all the steps of your task before you start.]
Answer index: 1
Subject: Language Arts
Skill: show-character-emotions-and-traits
Description: You are writing a story about Devon, who hears some bad news. Choose the best description to show that Devon is **disappointed**.
Choices: [(a) Devon swallowed hard and tried to force herself to smile., (b) "Okay," Devon said, continuing to chop carrots.]
Answer index: 0
@@
Subject: Language Arts
Skill: identify-text-structures
Description: Read the text.
@@
Although there are many kinds of writing, most people divide texts into two different types: fiction and nonfiction. Fiction means stories that are made up. "Charlotte’s Web" is fiction because, of course, spiders and pigs don’t really talk. Nonfiction, however, gives you facts about the real world. You can find nonfiction books about history, science, real people, animals, and many other topics. Nonfiction also includes newspapers and textbooks, while fiction includes stories, plays, and novels. Reading both fiction and nonfiction can be fun and educational.
Which text structure does the text use?
Choices: [(a) cause-effect, (b) compare-contrast]
Answer index: 1
Subject: Language Arts
Skill: which-sentence-matches-the-definition
Description: Which sentence uses this meaning of **concrete**?
*concrete** "adjective" specific and real
@@
Choices: [(a) His fingerprints were **concrete** proof that he had opened the safe., (b) We can’t walk on the sidewalk because the **concrete** is still drying.]
Answer index: 0
Subject: Language Arts
Skill: determine-the-meaning-of-idioms-from-context-set-1
Description: What is the meaning of **on the tip of one’s tongue**?
I can’t quite remember the name of that magician we saw perform in Las Vegas, but it is **on the tip of my tongue**.
Choices: [(a) almost remembered, (b) a certainty]
Answer index: 0
Subject: Language Arts
Skill: use-guide-words
@@
Description: Which word would you find on a dictionary page with the following guide words?
*skate** - **squire**
Choices: [(a) spare, (b) station]
Answer index: 0
Subject: Language Arts
Skill: use-context-to-identify-the-meaning-of-a-word
Description: Read the passage and then answer the question.
@@
The Great Pyramid of Giza is the largest of the ancient Egyptian pyramids. When completed around 2560 BCE, it stood 481 feet high, as tall as a modern 44-story building. The Great Pyramid was the world’s tallest building for more than three thousand years. More than two million huge stone blocks were used to construct it. The builders of the pyramid positioned it **meticulously** so that it perfectly faces true north. Scientists aren’t sure how the ancient Egyptians knew how to place the base. Experts also disagree about how those huge blocks were moved into place. They don’t even fully understand the design of the inside of the pyramid. More work must be done to solve these mysteries.
What is the meaning of **meticulously** as used in the passage?
Choices: [(a) with many people pushing, (b) in a great hurry, (c) happily, without any hesitation, (d) carefully, with attention to detail]
Answer index: 3
Subject: Language Arts
Skill: punctuating-dialogue
Description: Which sentence is correct?
@@
Choices: [(a) "Please hire a carpenter to finish our basement", Mrs. Huffman urged., (b) "Please hire a carpenter to finish our basement," Mrs. Huffman urged.]
Answer index: 1
Subject: Language Arts
Skill: determine-the-meanings-of-similes-and-metaphors
Description: What is the meaning of the simile?
Linda’s typing was like a tap dancer performing onstage.
Choices: [(a) Linda made rapid tapping noises., (b) Linda made her performance on the stage.]
Answer index: 0
Subject: Language Arts
Skill: use-key-details-to-determine-the-main-idea
Description: Look at these details from a paragraph about the same topic:
By taking music lessons, students learn to read sheet music.
Through music lessons, students can learn about music from other countries.
Choose the main, or central, idea that ties all the details together.
Choices: [(a) Music lessons introduce students to many types of music., (b) Music lessons teach more than just how to play an instrument., (c) Piano lessons are one of the most popular types of music lesson.]
Answer index: 1
Subject: Language Arts
Skill: choose-between-adjectives-and-adverbs
Description: Choose the word that best describes the adjective in bold.
As a wedding gift, Elena gave the couple a quilt that was ___ **sewn** by hand.
Choices: [(a) full, (b) fully]
Answer index: 1
Subject: Language Arts
Skill: read-realistic-fiction
Description: Read the story.
title: The Artist’s Secret
Destiny smiled. Then she reached for her palette, so she could start on C.
Based on paragraphs 1 and 2, how does Destiny probably feel about the art project at the beginning of the story?
Choices: [(a) She cares about it turning out just right., (b) She wants to finish it as quickly as possible., (c) She views it as a competition that she is winning., (d) She is annoyed about having to mix paint colors.]
Answer index: 0
Subject: Language Arts
Skill: use-the-meanings-of-words-as-clues-to-the-meanings-of-greek-and-latin-roots
Description: A biography is a written account of someone’s life. What does the root **graph** mean?
Choices: [(a) look at or observe, (b) write or draw, (c) measure]
Answer index: 1
Subject: Language Arts
Skill: use-coordinating-conjunctions
Description: Complete the sentence with the best conjunction.
Pam wants to dye her hair purple ___ red.
Choices: [(a) or, (b) but]
Answer index: 0
Subject: Language Arts
Skill: identify-similes-and-metaphors
Description: Does the sentence use a simile or a metaphor?
Mrs. Boyer is kind, and her heart is gold.
Choices: [(a) simile, (b) metaphor]
Answer index: 1
Subject: Language Arts
Skill: use-greek-and-latin-roots-as-clues-to-the-meanings-of-words
Description: The root **vis** means "see or look at". What does the word **visual** mean?
Choices: [(a) relating to hearing or sound, (b) relating to seeing or sight, (c) relating to flavor or taste]
Answer index: 1
Subject: Language Arts
Skill: interpret-the-meaning-of-an-allusion-from-its-source
Description: Read the source of the allusion.
@@
In the fairy tale, Cinderella is a poor young woman who is treated badly. However, she later marries a prince and becomes a loved and respected princess.
What is the meaning of the allusion in the sentence below?
Choices: [(a) an amazing improvement, (b) a journey without troubles]
Answer index: 0
Subject: Language Arts
Skill: positive-and-negative-connotation
Description: Which phrase has a more **negative** connotation?
Choices: [(a) a warm room, (b) a stuffy room]
Answer index: 1